On January 1, 2019, Shay Company issues $430,000 of 8%, 15-year bonds. The bonds sell for $417,100. Six years later, on January 1, 2025, Shay retires these bonds by buying them on the open market for $450,425. All interest is accounted for and paid through December 31, 2024, the day before the purchase. The straight-line method is used to amortize any bond discount.

1. What is the amount of the discount on the bonds at issuance?
2. How much amortization of the discount is recorded on the bonds for the entire period from January 1, 2019, through December 31, 2024?
3. What is the carrying (book) value of the bonds as of the close of business on December 31, 2024?
4. Prepare the journal entry to record the bond retirement.
Answer #1

1) Discount on bond issuance = 430000-417100 = 12900

2) Amortization of discount from January 1,2019 to Dec 31,2024 = 12900/15*6 = 5160

3) Carrying value on December 31,2024 = 417100+5160 = 422260

4) Journal entry

Date account and explanation Debit Credit
Jan 1,2025 Bonds payable 430000
Loss on retirement of bonds 28165
Discount on bonds payable 7740
Cash 450425